# Variational Bayes Gaussian Splatting

### v1

@@ 5 Discussion and Conclusion
In this paper, we introduced a novel approach to optimizing Gaussian splats using variational Bayes. Our approach achieves comparable performance to backpropagation-based methods on 2D and 3D datasets, while offering key advantages for scenarios involving continual learning from streaming data.
One limitation of VBGS compared to 3D Gaussian Splatting (3DGS) is its reliance on RGBD data, as opposed to optimizing directly on RGB projections. For many use cases we envision, such as robot navigation, depth information is often readily available from stereo vision, lidar or other sensing technologies. Furthermore, most 3DGS approaches still rely on pre-computed camera parameters obtained from structure-from-motion algorithms (Özyeşil et al., 2017), which implicitly also provide depth information through triangulation. One could also use a pretrained neural network that predicts depth from monocular RGB data (Yang et al., 2024), as evidenced by Fu et al. (2024).
Additionally, 3DGS dynamically adjusts the model size by growing and shrinking as needed, a feature that we have not yet incorporated. In future work, we aim to explore principled approaches for dynamic model sizing, leveraging model evidence (Friston et al., 2023) to guide this process.
While VBGS requires only a single update step per observation, this step is computationally more expensive than a single gradient descent iteration in backpropagation. For extremely large datasets that do not fit into memory, VBGS may need to process data in distinct update steps, where stochastic gradient descent (SGD) might offer better efficiency. However, in streaming or partial data settings, VBGS excels by allowing updates without replay buffers, opening avenues for active data selection, a potential future research direction. Investigating active learning mechanisms, particularly in robotic SLAM or autonomous systems, could optimize data usage, ensuring that only the most informative observations are processed.
@@
6 Code availability
The code for using VBGS is available in the vbgs repository, which can be found at the following link: https://github.com/VersesTech/vbgs.
7 Acknowledgments and Disclosure of Funding
The authors would like to thank the members of the VERSES AI Research lab for critical discussions and feedback that improved the quality of this work, with special thanks to Jeff Beck, Tommaso Salvatori, and Conor Heins.
@@ A.1 Prior parameters
Table 2: Parameters of the conjugate priors over likelihood parameters. Some values are a function of the number of components, indicated by nc, $I$ represents the identity matrix of size the multivariate dimension $D$. Parameters are in the canonical form of the corresponding distribution.
@@
$V_{{s},k}$ $2.25\cdot 10^{4}\cdot\text{nc}\cdot I$ $2.25\cdot 10^{6}\cdot\text{nc}\cdot I$
@@ A.2 Initial parameters
Table 3: Parameters of the initial approximate posteriors over likelihood parameters. Some values are a function of the number of components, indicated by nc, $I$ represents the identity matrix of size the multivariate dimension $D$. Parameters are in the canonical form of the corresponding distribution.
@@
$V_{{s},k}$ $2.25\cdot 10^{4}\cdot\text{nc}\cdot I$ $2.25\cdot 10^{6}\cdot\text{nc}\cdot I$
@@ D.2 Growing and shrinking of 3D Gaussian Splatting
Table 5: Dynamic 3DGS. The first row shows reconstruction performance, measured in PSNR (dB), evaluated on the validation set of the eight objects of the Blender set. The second row for each model shows the resulting amount of components after optimization.
@@
Gradient (Data ${}^{\text{(Data Init)}}$ 26.73 ±1.15 20.65 ±0.71 22.93 ±1.09 27.92 ±0.84 26.94 ±1.10 17.02 ±1.48 28.10 ±0.81 25.91 ±1.62
@@
Gradient (Random ${}^{\text{(Random Init)}}$ 26.72 ±1.16 20.65 ±0.69 22.94 ±1.10 28.06 ±0.80 26.98 ±1.11 17.03 ±1.48 28.26 ±0.79 25.81 ±1.59
@@ D.3 Novel view prediction performance on Habitat rooms
The results for the various model configurations with 100K components on the rooms data are shown in Table 6. We observe that the performance of all models lies within the same range, with the gradient descent approach yielding better performance when initialized on the data. However, VBGS performs better when initialized randomly. Using component reassignments also pushes the performance of Random Init close to Data Init while outperforming it for “Apartment 1”.
@@
VBGS (Data ${}^{\text{(Data Init)}}$ $19.83$4.56 $\mathbf{18.66}$2.02 $22.92$3.33
VBGS (Random Init - ${}^{\text{(Random Init - Reassign)}}$ $19.78$4.49 $17.93$2.52 $22.98$3.72
VBGS (Random ${}^{\text{(Random Init)}}$ $18.98$3.89 $15.62$2.27 $20.38$4.53
Gradient (Data ${}^{\text{(Data Init)}}$ $\mathbf{20.97}$4.29 $18.45$1.89 $\mathbf{24.96}$6.17
Gradient (Random ${}^{\text{(Random Init)}}$ $20.55$4.02 $14.30$4.36 $19.18$5.32

### v2

@@ 5 Discussion and Conclusion
In this paper, we introduced a novel approach to optimizing Gaussian splats using variational Bayes. Our approach achieves comparable performance to backpropagation-based methods on 2D and 3D datasets, while offering key advantages for scenarios with continual learning on streaming data.
One limitation of VBGS compared to 3D Gaussian Splatting (3DGS) is its reliance on RGBD data, as opposed to optimizing directly on RGB projections. For many use cases we envision, such as robot navigation, depth information is often readily available from stereo vision, lidar or other sensing technologies. Furthermore, most 3DGS approaches still rely on pre-computed camera parameters obtained from structure-from-motion algorithms (Özyeşil et al., 2017), which implicitly also provide depth information through triangulation. One could also use a pretrained neural network that predicts depth from monocular RGB data (Yang et al., 2024b), as evidenced by Fu et al. (2024) (please see Appendix D.5 for an experiment on this approach using VBGS).
As VBGS represents a distribution over the parameters of a Gaussian Mixture Model (GMM) rather than optimizing for the maximum likelihood solution directly, this increases the memory requirements by a factor of 2 compared to the gradient baseline. This richer parameterization allows training to be reformulated as coordinate ascent variational inference (CAVI). A single VBGS update step was found to be faster compared to the accumulated gradient update steps to reach the same performance, given enough memory. However, in large-scale 3D datasets, batched processing becomes necessary, losing this competitive advantage. For a detailed discussion, please see Appendix E.
However, in streaming or partial data settings, VBGS excels by allowing updates without replay buffers, opening avenues for active data selection, a potential future research direction. Investigating active learning mechanisms, particularly in robotic SLAM or autonomous systems, could optimize data usage, ensuring that only the most informative observations are processed.
@@ A.1 Prior parameters
Table 2: Parameters of the conjugate priors over likelihood parameters. Some values are a function of the number of components, indicated by nc, II represents the identity matrix of size the multivariate dimension $D$. Parameters are in the canonical form of the corresponding distribution.
@@
$V_{{s},k}$ $25\cdot 10^{4}\cdot\text{nc}\cdot I 2.2$$25\cdot 10^{6}\cdot\text{nc}\cdot I$
@@ A.2 Initial parameters
Table 3: Parameters of the initial approximate posteriors over likelihood parameters. Some values are a function of the number of components, indicated by nc, II represents the identity matrix of size the multivariate dimension $D$. Parameters are in the canonical form of the corresponding distribution.
@@
$V_{{s},k}$ $25\cdot 10^{4}\cdot\text{nc}\cdot I 2.2$$25\cdot 10^{6}\cdot\text{nc}\cdot I$
@@ D.2 Growing and shrinking of 3D Gaussian Splatting
Table 5: Dynamic 3DGS. The first row shows reconstruction performance, measured in PSNR (dB), evaluated on the validation set of the eight objects of the Blender set. The second row for each model shows the resulting amount of components after optimization. The top rows show the performance of dynamically growing 3DGS, while the bottom is the performance of a VBGS initialzied on the equivalent amount of components.
@@
Gradient $^{\text{(Data Init)}}$ 26.73 ±1.15 20.65 ±0.71 22.93 ±1.09 27.92 ±0.84 26.94 ±1.10 17.02 ±1.48 28.10 ±0.81 25.91 ±1.62
@@
Gradient $^{\text{(Random Init)}}$ 26.72 ±1.16 20.65 ±0.69 22.94 ±1.10 28.06 ±0.80 26.98 ±1.11 17.03 ±1.48 28.26 ±0.79 25.81 ±1.59
@@
VBGS $^{\text{(Data Init)}}$ $22.66$ ±0.93 $19.37$ ±0.47 $22.43$ ±0.77 $25.13$ ±0.99 $23.02$ ±0.84 $21.12$ ±1.57 $25.18$ ±0.63 $22.51$ ±0.86
281K $^{\text{/445K}}$ 242K $^{\text{/384K}}$ 178K $^{\text{/282K}}$  [^1] K $^{\text{/177K}}$ 214K $^{\text{/339K}}$ 33K $^{\text{/53K}}$ 114K $^{\text{/182K}}$ 165K $^{\text{/262K}}$
@@
Additionally, we initialize the VBGS approach on the amount of components 3DGS converges on. We observe that the performance achieved by VBGS is similar to the 100K model (see Table 1). Note that the objective of VBGS trades off accuracy with complexity, and hence a large amount of components remain unused.
We also evaluate on a dataset of 3D rooms from the Habitat test suite (“Van Gogh Room”, “Apartment 1” and “Skokloster Castle”) [Savva et al., 2019]. The model parameters are inferred using 200 randomly sampled views from each room using the preprocessing pipeline of [Wang et al., 2023]. On top of the models from the previous sections, we now also evaluate the model with component reassignments as detailed in Section 3.4.
@@ D.3 Novel view prediction performance on Habitat rooms
VBGS $^{\text{(Data Init)}}$ $19.83$ ±4.56 $\mathbf{18.66}$ ±2.02 $22.92$ ±3.33
VBGS $^{\text{(Random Init - Reassign)}}$ $19.78$ ±4.49 $17.93$ ±2.52 $22.98$ ±3.72
VBGS $^{\text{(Random Init)}}$ $18.98$ ±3.89 $15.62$ ±2.27 $20.38$ ±4.53
Gradient $^{\text{(Data Init)}}$ $\mathbf{20.97}$ ±4.29 $18.45$ ±1.89 $\mathbf{24.96}$ ±6.17
Gradient $^{\text{(Random Init)}}$ $20.55$ ±4.02 $14.30$ ±4.36 $19.18$ ±5.32
The results for the various model configurations with 100K components on the Habitat rooms data is shown in Table 7. We observe that for the random initial data condition (Random Init), VBGS performs better than Gradient for the “Skokloster Castle” and “Apartment 1”, while Gradient performs better for the “Van Gogh Room”. Using component reassignments, the VBGS performance can be improved even more. In the case of initialization on data (Data Init), the Gradient approach performs better on the “Van Gogh Room” and “Apartment 1”.
D.4 Continual Learning on the Replica Dataset
Figure 8 shows additional qualitative results of VBGS using continual learning with reassign on the Replica dataset.
Figure 8: Qualitative Results on Replica: Qualitative results of novel view prediction as a function of time for the five evaluated rooms of the Replica dataset.
D.5 Depth-free VBGS
A limitation of this approach compared to traditional 3D Gaussian Splatting is that it requires depth information to be present for every frame. In this section, we run an ablation where we do not have access to depth information, but instead use a pretrained neural network (Depth-Anything-V2-Large [Yang et al., 2024a], trained on the Hypersim dataset [Roberts et al., 2021]) to estimate the metric depth of each frame.
We test this approach on “Room 0” of the Replica dataset [Straub et al., 2019], using the same train and evaluation sequence used in Appendix D.4.
As we evaluate the model on a different dataset than it was trained on, we need to calibrate the depth estimation network. We do this using ten ground truth depth frames for which we estimate the scale using linear regression without intercept, and find that for this dataset the output of the network should be scaled with a factor of $0.7$.
Qualitative examples of reconstruction performance using the predicted depth model are shown in Figure 9. Comparing these to the results in Figure 8, we observe that the inaccuracies in the predicted depth result in larger representations of the objects (i.e. the chairs in frame 50 are now entirely covered by the flowers).
Table 7: Performance for novel view prediction on Replica Room 0 with estimated depth. This compares VBGS with 100k components using random initialization and reassignments for both ground truth depth and estimated depth using DepthAnythingV2 [Yang et al., 2024a]. Values are reported as PSNR (dB) ($\mu\pm\sigma$), for which the statistics are computed over the 100 validation frames.
Room 0
VBGS $^{\text{(Random Init - Reassign)}}$ $20.46$ ±1.69
VBGS $^{\text{(Random Init - Reassign)}}$ $^{\text{Depth Estimated}}$ $16.26$ ±0.98
Figure 9: Continual learning results for the Replica rooms dataset with estimated depth. (a) Qualitative results of novel view prediction as a function of time for “Room 0” where the depth information is inferred using DepthAnythingV2 [Yang et al., 2024a]. (b) Evolution of the reconstruction performance, measured as PSNR (dB), after feeding in consecutive images of the room in the case where depth information is available, as well as were the estimated depth information is used.
D.6 Robot Applications
We conducted a preliminary experiment using a mobile robot in the Habitat simulator. The robot observes RGBD and its pose at every frame and updates the model at each step. This is visualized in Figure 10. The figure shows a top-down projection of the means of the Gaussians. The structure of the environment is captured without forgetting the parts observed at the beginning of the sequence. This could, for example, be used for autonomous navigation and collision avoidance.
Additionally, we fitted a VBGS on the TUM dataset [Engelhard et al., 2011], a large-scale real-world dataset commonly used for evaluating SLAM algorithms. We train on 225 frames sampled equidistant in time across the whole sequence. Qualitative reconstruction results from the validation set are shown in Figure 10. It can be observed that the model can capture the structure of both sides of the table.
Figure 10: Robot scenarios using VBGS: (a) Continual mapping of an environment in the Habitat simulator [Savva et al., 2019] using a VBGS with 10K components with reassign (fraction=$0.05$). Shown is a top-down representation of the colored means of VBGS, where components with a z-value larger than 2m (the ceiling) are filtered out. The robot trajectory is marked in red, and the shaded gray background marks the ground truth navigable area. (b) VBGS (100K components, random init with reassign) using continual data on the TUM [Engelhard et al., 2011] dataset commonly used to evaluate SLAM algorithms.
Appendix E Computational cost
Instead of optimizing for the maximum likelihood parameters of a Gaussian Mixture Model directly, VBGS represents a distribution over these parameters. This means that instead of representing each Gaussian component with its color, mean, scale, rotation and opacity (14 parameters per component) as is done in the gradient baseline, we represent the parameters of the Normal-Inverse-Wishart likelihood for color, position and Dirichlet prior (29 parameters per component), which increases the memory requirements by a factor of 2 for the same number of Gaussians.
This representation allows us to turn “training” into a coordinate ascent variational inference (CAVI) algorithm, which can be decomposed in two steps: computing model assignments which has complexity O(nm) with n the number of data points and m the number of Gaussian components, and computing the parameter updates with complexity O(n). Note that we can parallelize these computations over both the number of components as well as data on a GPU.
We conducted an experiment where we compare the computational cost of a single update step with the Gradient approach. Specifically, we first execute a single VBGS update, and then train the gradient-based Gaussian splat until it reaches the same level of reconstruction performance as VBGS (measured in PSNR). The difference in variance between VBGS and the gradient approach can be attributed to the fact that VBGS has a fixed number of steps (1) whereas for the gradient approach this is dependent on the reconstruction quality.
Given that all the data fits in memory, VBGS is significantly faster than the gradient baseline. However, for the 3D experiments this is no longer the case and we need to resort to batched processing of the observed data. An interesting venue for future work would be to actively sample data points and only compute the update for these points.
All experiments were conducted on an NVIDIA GeForce 4090 GTX with 24GB of memory.
Figure 11: Computation time for VBGS and Gradient approach. Values are measured as wall clock time on an NVIDIA RTX 4090 in seconds over the validation frames of the ImageNet dataset. Both models have 1K components.